Skip to content
Textbook Guides

Solution Manual For Introduction To Mathematical Statistics, 7th Edition

Gain deeper insight into your textbook problems with Solution Manual For Introduction To Mathematical Statistics, 7th Edition, featuring well-explained solutions.

Preview 16 of 109 pages

See the first 16 pages exactly as they appear in the document.

  1. Page 1
    Read page 1 as text

    S OLUTIONS M ANUAL I NTRODUCTION TO M ATHEMATICAL S TATISTICS S EVENTH E DITION Robert Hogg University of Iowa Joseph McKean Western Michigan University Allen Craig

  2. Page 2
  3. Page 3
    Read page 3 as text

    Contents 1 Probability and Distributions 1 2 Multivariate Distributions 11 3 Some Special Distributions 19 4 Some Elementary Statistical Inferences 31 5 Consistency and Limiting Distributions 49 6 Maximum Likelihood Methods 53 7 Sufficiency 65 8 Optimal Tests of Hypotheses 77 9 Inferences about Normal Models 83 10 Nonparametric and Robust Statistics 93 11 Bayesian Statistics 103 iii

  4. Page 4
    Read page 4 as text

    Chapter 1 Probability and Distributions 1.2.1 Part (c): C 1 ∩ C 2 = { ( x, y ) : 1 < x < 2 , 1 < y < 2 } . 1.2.3 C 1 ∩ C 2 = { mary,mray } . 1.2.6 C k = { x : 1 /k ≤ x ≤ 1 − (1 /k ) } . 1.2.7 C k = { ( x, y ) : 0 ≤ x ≤ 1 /k, 0 ≤ y ≤ 1 /k } . 1.2.8 lim k →∞ C k = { x : 0 < x < 3 } . Note: neither the number 0 nor the number 3 is in any of the sets C k , k = 1 , 2 , 3 , . . . 1.2.9 Part (b): lim k →∞ C k = φ , because no point is in all the sets C k , k = 1 , 2 , 3 , . . . 1.2.11 Because f ( x ) = 0 when 1 ≤ x < 10, Q ( C 3 ) = ∫ 10 0 f ( x ) dx = ∫ 1 0 6 x (1 − x ) dx = 1 . 1.2.13 Part (c): Draw the region C carefully, noting that x < 2 / 3 because 3 x/ 2 < 1. Thus Q ( C ) = ∫ 2 / 3 0 [ ∫ 3 x/ 2 x/ 2 dy ] dx = ∫ 2 / 3 0 x dx = 2 / 9 . 1.2.16 Note that 25 = Q ( C ) = Q ( C 1 ) + Q ( C 2 ) − Q ( C 1 ∩ C 2 ) = 19 + 16 − Q ( C 1 ∩ C 2 ) . Hence, Q ( C 1 ∩ C 2 ) = 10. 1.2.17 By studying a Venn diagram with 3 intersecting sets, it should be true that 11 ≥ 8 + 6 + 5 − 3 − 2 − 1 = 13 . It is not, and the accuracy of the report should be questioned. 1

  5. Page 5
    Read page 5 as text

    2 Probability and Distributions 1.3.3 P ( C ) = 1 2 + 1 4 + 1 8 + · · · = 1 / 2 1 − (1 / 2) = 1 . 1.3.6 P ( C ) = ∫ ∞ −∞ e −| x | dx = ∫ 0 −∞ e x dx + ∫ ∞ 0 e − x dx = 2 6 = 1 . We must multiply by 1 / 2. 1.3.8 P ( C c 1 ∪ C c 2 ) = P [( C 1 ∩ C 2 ) c ] = P ( C ) = 1 , because C 1 ∩ C 2 = φ and φ c = C . 1.3.11 The probability that he does not win a prize is ( 990 5 ) / ( 1000 5 ) . 1.3.13 Part (a): We must have 3 even or one even, 2 odd to have an even sum. Hence, the answer is ( 10 3 )( 10 0 ) ( 20 3 ) + ( 10 1 )( 10 2 ) ( 20 3 ) . 1.3.14 There are 5 mutual exclusive ways this can happen: two “ones”, two “twos”, two “threes”, two “reds”, two “blues.” The sum of the corresponding proba- bilities is ( 2 2 )( 6 0 ) + ( 2 2 )( 6 0 ) + ( 2 2 )( 6 0 ) + ( 5 2 )( 3 0 ) + ( 3 2 )( 5 0 ) ( 8 2 ) . 1.3.15 (a) 1 − ( 48 5 )( 2 0 ) ( 50 5 ) (b) 1 − ( 48 n )( 2 0 ) ( 50 n ) ≥ 1 2 , Solve for n . 1.3.20 Choose an integer n 0 > max { a − 1 , (1 − a ) − 1 } . Then { a } = ∩ ∞ n = n 0 ( a − 1 n , a + 1 n ) . Hence by (1.3.10), P ( { a } ) = lim n →∞ P [( a − 1 n , a + 1 n )] = 2 n = 0 . 1.4.2 P [( C 1 ∩ C 2 ∩ C 3 ) ∩ C 4 ] = P [ C 4 | C 1 ∩ C 2 ∩ C 3 ] P ( C 1 ∩ C 2 ∩ C 3 ) , and so forth. That is, write the last factor as P [( C 1 ∩ C 2 ) ∩ C 3 ] = P [ C 3 | C 1 ∩ C 2 ] P ( C 1 ∩ C 2 ) .

  6. Page 6
    Read page 6 as text

    3 1.4.5 [( 4 3 )( 48 10 ) + ( 4 4 )( 48 9 )] / ( 52 13 ) [( 4 2 )( 48 11 ) + ( 4 3 )( 48 10 ) + ( 4 4 )( 48 9 )] / ( 52 13 ) . 1.4.10 P ( C 1 | C ) = (2 / 3)(3 / 10) (2 / 3)(3 / 10) + (1 / 3)(8 / 10) = 3 7 < 2 3 = P ( C 1 ) . 1.4.12 Part (c): P ( C 1 ∪ C c 2 ) = 1 − P [( C 1 ∪ C c 2 ) c ] = 1 − P ( C ∗ 1 ∩ C 2 ) = 1 − (0 . 4)(0 . 3) = 0 . 88 . 1.4.14 Part (d): 1 − (0 . 3) 2 (0 . 1)(0 . 6) . 1.4.16 1 − P ( T T ) = 1 − (1 / 2)(1 / 2) = 3 / 4, assuming independence and that H and T are equilikely. 1.4.19 Let C be the complement of the event; i.e., C equals at most 3 draws to get the first spade. (a) P ( C ) = 1 4 + 3 4 1 4 + ( 3 4 ) 2 1 4 . (b) P ( C ) = 1 4 + 13 51 39 52 + 13 50 38 51 39 52 . 1.4.22 The probability that A wins is ∑ ∞ n =0 ( 5 6 4 6 ) n 1 6 = 3 8 . 1.4.27 Let Y denote the bulb is yellow and let T 1 and T 2 denote bags of the first and second types, respectively. (a) P ( Y ) = P ( Y | T 1 ) P ( T 1 ) + P ( Y | T 2 ) P ( T 2 ) = 20 25 . 6 + 10 25 . 4 . (b) P ( T 1 | Y ) = P ( Y | T 1 ) P ( T 1 ) P ( Y ) . 1.4.30 Suppose without loss of generality that the prize is behind curtain 1. Con- dition on the event that the contestant switches. If the contestant chooses curtain 2 then she wins, (In this case Monte cannot choose curtain 1, so he must choose curtain 3 and, hence, the contestant switches to curtain 1). Like- wise, in the case the contestant chooses curtain 3. If the contestant chooses curtain 1, she loses. Therefore the conditional probability that she wins is 2 3 . 1.4.31 (1) The probability is 1 − ( 5 6 ) 4 . (2) The probability is 1 − [ ( 5 6 ) 2 + 10 36 ] 24 .

  7. Page 7
    Read page 7 as text

    4 Probability and Distributions 1.5.2 Part (a): c [(2 / 3) + (2 / 3) 2 + (2 / 3) 3 + · · · ] = c (2 / 3) 1 − (2 / 3) = 2 c = 1 , so c = 1 / 2. 1.5.5 Part (a): p ( x ) = { ( 13 x )( 39 5 − x ) ( 52 5 ) x = 0 , 1 , . . . , 5 0 elsewhere . 1.5.9 Part (b): 50 ∑ x =1 x/ 5050 = 50(51) 2(5050) = 51 202 . 1.5.10 For Part (c): Let C n = { X ≤ n } . Then C n ⊂ C n +1 and ∪ n C n = R . Hence, lim n →∞ F ( n ) = 1. Let ǫ > 0 be given. Choose n 0 such that n ≥ n 0 implies 1 − F ( n ) < ǫ . Then if x ≥ n 0 , 1 − F ( x ) ≤ 1 − F ( n 0 ) < ǫ . 1.6.2 Part (a): p ( x ) = ( 9 x − 1 ) ( 10 x − 1 ) 1 11 − x = 1 10 , x = 1 , 2 , . . . 10 . 1.6.3 (a) p ( x ) = ( 5 6 ) x − 1 ( 1 6 ) , x = 1 , 2 , 3 , . . . (b) ∞ ∑ x =1 ( 5 6 ) x − 1 ( 1 6 ) = 1 / 6 1 − (25 / 36) = 6 11 . 1.6.8 D y = { 1 , 2 3 , 3 3 , . . . } . The pmf of Y is p ( y ) = ( 1 2 ) y 1 / 3 , y ∈ D y . 1.7.1 If √ x < 10 then F ( x ) = P [ X ( c ) = c 2 ≤ x ] = P ( c ≤ √ x ) = ∫ √ x 0 1 10 dz = √ x 10 . Thus f ( x ) = F ′ ( x ) = { 1 20 √ x 0 < x < 100 0 elsewhere . 1.7.2 C 2 ⊂ C c 1 ⇒ P ( C 2 ) ≤ P ( C c 1 ) = 1 − (3 / 8) = 5 / 8 .

  8. Page 8
    Read page 8 as text

    5 1.7.4 Among other characteristics, ∫ ∞ −∞ 1 π (1 + x 2 ) dx = 1 π arctan x ∣ ∣ ∣ ∣ ∣ ∣ ∞ −∞ = 1 π [ π 2 − ( − π 2 )] = 1 . 1.7.6 Part (b): P ( X 2 < 9) = P ( − 3 < X < 3) = ∫ 3 − 2 x + 2 19 dx = 1 18 [ x 2 2 + 2 x ] 3 − 2 = 1 18 [ 21 2 − ( − 2) ] = 25 36 . 1.7.8 Part (c): f ′ ( x ) = 1 2 2 xe − x = 0; hence, x = 2 is the mode because it maximizes f ( x ). 1.7.9 Part (b): ∫ m 0 3 x 2 dx = 1 2 ; hence, m 3 = 2 − 1 and m = (1 / 2) 1 / 3 . 1.7.10 ∫ ξ 0 . 2 0 4 x 3 dx = 0 . 2 : hence, ξ 4 0 . 2 = 0 . 2 and ξ 0 . 2 = 0 . 2 1 / 4 . 1.7.13 x = 1 is the mode because for 0 < x < ∞ because f ( x ) = F ′ ( x ) = e − x − e − x + xe − x = xe − x f ′ ( x ) = − xe − x + e − x = 0 , and f ′ (1) = 0. 1.7.16 Since ∆ > 0 X > z ⇒ Y = X + ∆ > z. Hence, P ( X > z ) ≤ P ( Y > z ). 1.7.19 Since f ( x ) is symmetric about 0, ξ . 25 < 0. So we need to solve, ∫ ξ . 25 − 2 ( − x 4 ) dx = . 25 . The solution is ξ . 25 = − √ 2.

  9. Page 9
    Read page 9 as text

    6 Probability and Distributions 1.7.20 For 0 < y < 27, x = y 1 / 3 , dx dy = 1 3 y − 2 / 3 g ( y ) = = 1 3 y 2 / 3 y 2 / 3 9 = 1 27 . 1.7.22 f ( x ) = 1 π , − π 2 < x < π 2 . x = arctan y, dx dy = 1 1 + y 2 , −∞ < y < ∞ . g ( y ) = 1 π 1 1 + y 2 , −∞ < y < ∞ . 1.7.23 G ( y ) = P ( − 2 log X 4 ≤ y ) = P ( X ≥ e − y/ 8 ) = ∫ 1 e − y/ 8 4 x 3 dx = 1 − e − y/ 2 , 0 < y < ∞ g ( y ) = G ′ ( y ) = { e − y/ 2 0 < y < ∞ 0 elsewhere . 1.7.24 G ( y ) = P ( X 2 ≤ y ) = P ( − √ y ≤ X ≤ √ y ) =    ∫ √ y − √ y 1 3 dx = 2 √ y 3 0 ≤ y < 1 ∫ √ y − 1 1 3 dx = √ y 3 + 1 3 1 ≤ y < 4 g ( y ) =          1 3 √ y 0 ≤ y < 1 1 6 √ y 1 ≤ y < 4 0 elsewhere . 1.8.4 E (1 /X ) = 100 ∑ x =51 1 x 1 50 . The latter sum is bounded by the two integrals ∫ 101 51 1 x dx and ∫ 100 50 1 x dx. An appropriate approximation might be 1 50 ∫ 101 . 5 50 . 5 1 x dx = 1 50 (log 100 . 5 − log 50 . 5) .

  10. Page 10
    Read page 10 as text

    7 1.8.6 E [ X (1 − X )] = ∫ 1 0 x (1 − x )3 x 2 dx. 1.8.8 When 1 < y < ∞ G ( y ) = P (1 /X ≤ y ) = P ( X ≥ 1 /y ) = ∫ 1 1 /y 2 x dx = 1 − 1 y 2 g ( y ) = 2 y 3 E ( Y ) = ∫ ∞ 1 y 2 y 3 dy = 2 , which equals ∫ 1 0 (1 /x )2 x dx . 1.8.10 The expectation of X does not exist because E ( | X | ) = 2 π ∫ ∞ 0 x 1 + x 2 dx = 1 π ∫ ∞ 1 1 u du = ∞ , where the change of variable u = 1 + x 2 was used. 1.9.2 M ( t ) = ∞ ∑ x =1 ( e t 2 ) x = e t / 2 1 − ( e t / 2) , e t / 2 < 1 . Find E ( X ) = M ′ (0) and Var( X ) = M ′′ (0) − [ M ′ (0)] 2 . 1.9.4 0 ≤ var( X ) = E ( X 2 ) − [ E ( X )] 2 . 1.9.6 E [ ( X − μ σ ) 2 ] = 1 σ 2 σ 2 = 1 . 1.9.8 K ( b ) = E [( X − b ) 2 ] = E ( X 2 ) − 2 bE ( X ) + b 2 K ′ ( b ) = − 2 E ( X ) + 2 b = 0 ⇒ b = E ( X ) . 1.9.11 For a continuous type random variable, K ( t ) = ∫ ∞ −∞ t x f ( x ) dx. K ′ ( t ) = ∫ ∞ −∞ xt x − 1 f ( x ) dx ⇒ K ′ (1) = E ( X ) . K ′′ ( t ) = ∫ ∞ −∞ x ( x − 1) t x − 2 f ( x ) dx ⇒ K ′′ (1) = E [ X ( X 1 )]; and so forth.

  11. Page 11
    Read page 11 as text

    8 Probability and Distributions 1.9.12 3 = E ( X − 7) ⇒ E ( X ) = 10 = μ. 11 = E [( X − 7) 2 ] = E ( X 2 ) − 14 E ( X ) + 49 = E ( X 2 ) − 91 ⇒ E ( X 2 ) = 102 and var( X ) = 102 − 100 = 2 . 15 = E [( X − 7) 3 ] . Expand ( X − 7) 3 and continue. 1.9.16 E ( X ) = 0 ⇒ var( X ) = E ( X 2 ) = 2 p. E ( X 4 ) = 2 p ⇒ kurtosis = 2 p/ 4 p 2 = 1 / 2 p. 1.9.17 ψ ′ ( t ) = M ′ ( t ) /M ( t ) ⇒ ψ ′ (0) = M ′ (0) /M (0) = E ( X ) . ψ ′′ ( t ) = M ( t ) M ′′ ( t ) − M ′ ( t ) M ′ ( t ) [ M ( t ) 2 ] ⇒ ψ ′′ (0) = M (0) M ′′ (0) − M ′ (0) M ′ (0) [ M (0) 2 ] = M ′′ (0) − [ M ′ (0)] 2 = var( X ) . 1.9.19 M ( t ) = (1 − t ) − 3 = 1 + 3 t + 3 · 4 t 2 2! + 3 · 4 · 5 t 3 3! + · · · Considering the coefficient of t r /r !, we have E ( X r ) = 3 · 4 · 5 · · · ( r + 2) , r = 1 , 2 , 3 . . . . 1.9.20 Integrating the parts with u = 1 − F ( x ), dv = dx , we get { [1 − F ( x )] x } b 0 − ∫ b 0 x [ − f ( x )] dx = ∫ b 0 xf ( x ) dx = E ( X ) . 1.9.23 E ( X ) = ∫ 1 0 x 1 4 dx + 0 · 1 4 + 1 · 1 2 = 5 8 . E ( X 2 ) = ∫ 1 0 x 2 1 4 dx + 0 · 1 4 + 1 · 1 2 = 7 12 . var( X ) = 7 12 − ( 5 8 ) 2 = 37 192 . 1.9.24 E ( X ) = ∫ ∞ −∞ x [ c 1 f 1 ( x ) + · · · + c k f k ( x )] dx = k ∑ i =1 c i μ i = μ.

  12. Page 12
    Read page 12 as text

    9 Because ∫ ∞ −∞ ( x − μ ) 2 f i ( x ) dx = σ 2 i + ( μ i − μ ) 2 , we have E [( X − μ ) 2 ] = k ∑ i =1 c i [ σ 2 i + ( μ i − μ ) 2 ] . 1.10.2 μ = ∫ ∞ 0 xf ( x ) dx ≥ ∫ ∞ 2 μ 2 μf ( x ) dx = 2 μP ( X > 2 μ ) . Thus 1 2 ≥ P ( X > 2 μ ) . 1.10.4 If, in Theorem 1.10.2, we take u ( X ) = exp { tX } and c = exp { ta } , we have P (exp { tX } ≥ exp { ta } ] ≤ M ( t ) exp {− ta } . If t > 0, the events exp { tX } ≥ exp { ta } and X ≥ a are equivalent. If t < 0, the events exp { tX } ≥ exp { ta } and X ≤ a are equivalent. 1.10.5 We have P ( X ≥ 1) ≤ [1 − exp {− 2 t } ] / 2 t for all 0 < t < ∞ , and P ( X ≤ − 1) ≤ [exp { 2 t } − 1] / 2 t for all −∞ < t < 0. Each of these bounds has the limit 0 as t → ∞ and t → −∞ , respectively.

  13. Page 13
  14. Page 14
    Read page 14 as text

    Chapter 2 Multivariate Distributions 2.1.2 P ( A 5 ) = 7 8 − 4 8 − 3 8 + 2 8 = 2 8 . 2.1.5 ∫ ∞ 0 ∫ ∞ 0 [ 2 g ( √ x 2 1 + x 2 2 ) /π √ x 2 1 + x 2 2 ] dx 1 dx 2 = ∫ ∞ 0 ∫ π/ 2 0 [2 g ( ρ ) /πρ ] ρ dθdρ = ∫ ∞ 0 g ( ρ ) dρ = 1 . 2.1.6 G ( z ) = P ( X + Y ≤ z ) = ∫ z 0 ∫ z − x 0 e − x − y dydx = ∫ z 0 [1 − e − ( z − x ) ] e − x dx = 1 − e − z − ze − z . g ( z ) = G ′ ( z ) = { ze − z 0 < z < ∞ 0 elsewhere . 2.1.7 G ( z ) = P ( XY ≤ z ) = 1 − ∫ 1 z ∫ 1 z/x dydx = 1 − ∫ 1 z ( 1 − z x ) dx = z − z log z g ( z ) = G ′ ( z ) = { − log z 0 < z < 1 0 elsewhere . Why is − log z > 0? 11

  15. Page 15
    Read page 15 as text

    12 Multivariate Distributions 2.1.8 f ( x, y ) =    ( 13 x )( 13 y )( 26 13 − x − y ) ( 52 13 ) x ≥ 0 , y ≥ 0 , x + y ≤ 13 , x and y integers 0 elsewhere . 2.1.10 P ( X 1 + X 2 ≤ 1) = 15 ∫ 1 / 2 0 x 2 1 [∫ 1 − x 1 x 1 x 2 dx 2 ] dx 1 . 2.1.14 E [ e t 1 X 1 + t 2 X 2 ] = ∞ ∑ i =1 ∞ ∑ j =1 e t 1 i + t 2 j ( 1 2 ) i + j = ∞ ∑ i =1 ( e t 1 1 2 ) i ∞ ∑ j =1 ( e t 1 1 2 ) j = [ 1 1 − 2 − 1 e t 1 − 1 ] [ 1 1 − 2 − 1 e t 2 − 1 ] , provided t i < log 2, i = 1 , 2. 2.2.1 p ( y 1 , y 2 ) = { ( 2 3 ) y 2 ( 1 3 ) 2 − y 2 ( y 1 , y 2 ) = (0 , 0) , ( − 1 , 1) , (1 , 1) , (0 , 2) 0 elsewhere . 2.2.2 p ( y 1 , y 2 ) = { y 1 / 36 y 1 = y 2 , 2 y 2 , 3 y 2 ; y 2 = 1 , 2 , 3 0 elsewhere . y 1 1 2 3 4 6 9 p ( y 1 ) 1/36 4/36 6/36 4/36 12/36 9/36 2.2.4 The inverse transformation is given by x 1 = y 1 y 2 and x 2 = y 2 with Jacobian J = y 2 . By noting what the boundaries of the space S ( X 1 , X 2 ) map into, it follows that the space T ( Y 1 , Y 2 ) = { ( y 1 , y 2 ) : 0 < y i < 1 , i = 1 , 2 } . The pdf of ( Y 1 , Y 2 ) is f Y 1 ,Y 2 ( y 1 , y 2 ) = 8 y 1 y 3 2 . 2.2.5 The inverse transformation is x 1 = y 1 − y 2 and x 2 = y 2 with Jacobian J = 1. The space of ( Y 1 , Y 2 ) is T = { ( y 1 , y 2 ) : −∞ < y i < ∞ , i = 1 , 2 } . Thus the joint pdf of ( Y 1 , Y 2 ) is f Y 1 ,Y 2 ( y 1 , y 2 ) = f X 1 ,X 2 ( y 1 − y 2 , y 2 ) , which leads to formula (2.2.1).

  16. Page 16
    Read page 16 as text

    13 2.3.2 ( a ) c 1 ∫ x 2 0 x 1 /x 2 2 dx 1 = c 1 2 = 1 ⇒ c 1 = 2 and c 2 = 5 . ( b ) 10 x 1 x 2 2 , 0 < x 1 < x 2 < 1; zero elsewhere ( c ) ∫ 1 / 2 1 / 4 2 x 1 / (5 / 8) 2 dx = 64 25 ( 1 4 − 1 16 ) = 12 25 . ( d ) ∫ 1 / 2 1 / 4 ∫ 1 x 1 10 x 1 x 2 2 dx 2 dx 1 = ∫ 1 / 2 1 / 4 10 3 x 1 (1 − x 3 1 ) dx 1 = 135 512 . 2.3.3 f 2 ( x 2 ) = ∫ x 2 0 21 x 2 1 x 3 2 dx 1 = 7 x 6 2 , 0 < x 2 < 1 . f 1 | 2 ( x 1 | x 2 ) = 21 x 2 1 x 3 2 / 7 x 6 2 = 3 x 2 1 /x 3 2 , 0 < x 1 < x 2 . E ( X 1 | x 2 ) = ∫ x 2 0 x 1 (3 x 2 1 /x 3 2 ) dx 1 = 3 4 x 2 . G ( y ) = P ( 3 4 X 2 ≤ y ) = ∫ 4 y/ 3 0 7 x 6 2 dx 2 = ( 4 y 3 ) 7 , 0 < y < 3 4 g ( y ) = { 7 ( 4 3 ) 7 y 6 0 < y < 3 4 0 elsewhere . E ( Y ) = 7 8 3 4 = 21 32 . Var( Y ) = 7 1024 . E ( X 1 ) = 21 32 . Var( X 1 ) = 553 15360 > 7 1024 . 2.3.8 The marginal pdf of X is f X ( x ) = 2 ∫ ∞ x e − x e − y dy = 2 e − 2 x , 0 < x < ∞ . Hence, the conditional pdf of Y given X = x is f Y | X ( y | x ) = 2 e − x e − y 2 e − 2 x = e − ( y − x ) , 0 < x < y < ∞ , with conditional mean E ( Y | X = x ) = ∫ ∞ x ye − ( y − x ) dy = x + 1 , x > 0 .

93 more pages in the full document

Unlock it once and it stays in your library, ready to chat with or turn into flashcards.

Unlock the full document
See all

Study this document with CramX

Once it is in your library, every tool can work from it.